# Mapping Drivers of Greenness: Spatial Variable Selection for MODIS Vegetation Indices

### v1

@@ Abstract
keywords:
Spatially varying coefficients, Bayesian variable selection, group lasso, spatial statistics, vegetation indices
@@ 1 Introduction
The western United States is undergoing rapid and unprecedented ecological change, with vegetation increasingly exposed to severe drought stress and heightened wildfire risk [abatzoglou2016impact, williams2019observed, williams2020large]. These changes carry enormous economic costs: wildfires in the region are estimated to cost the U.S. economy \$394–\$893 billion annually, with damages exceeding \$90 billion during 2017–2021 alone. Accurately identifying which environmental factors govern vegetation health and productivity is therefore crucial for predicting ecosystem responses and designing effective adaptation strategies [running2004continuous, pettorelli2005using].
@@
This challenge is amplified by the region’s extreme environmental heterogeneity. The western United States spans from coastal rainforests receiving over 3000 mm of annual precipitation to interior deserts with less than 200 mm [diffenbaugh2015anthropogenic], leading to fundamentally different ecological constraints. At high elevations, temperature often limits productivity, whereas in arid lowlands, moisture availability is the dominant control [sun2024alternating, mukhtar2025elevation]. Effective vegetation monitoring in such a setting requires identifying which environmental variables exert substantively nonzero effects on vegetation across space, and determining where within the domain those effects are credibly nonzero.
@@
We address this problem using a predictive modeling framework for MODIS Enhanced Vegetation Index (EVI) data from the western United States, leveraging a rich suite of environmental predictors: spectral reflectance bands (red, near-infrared, blue, mid-infrared), ecosystem productivity and energy fluxes (Gross Primary Productivity, Latent Heat Flux), observation geometry (view zenith angle, sun zenith angle, relative azimuth angle), and land surface characteristics (land cover type). Standard linear regression models assume that coefficients are spatially constant, potentially obscuring important geographic variation in these relationships [fotheringham2002geographically]. At the same time, not all candidate drivers have effects that are credibly different from zero across the region. Systematically screening predictors for non-negligible effects improves interpretability and directs scientific focus toward drivers that truly matter. Predictors with spatially extensive, credibly nonzero effects motivate spatially targeted interpretation and monitoring efforts, whereas predictors with negligible influence can often be excluded or modeled more simply, reducing monitoring and modeling complexity without sacrificing predictive accuracy.
@@
Spatially varying coefficient (SVC) models are well suited to this problem because they allow regression relationships to change smoothly across space, producing coefficient surfaces that capture geographic variation in environmental effects [brunsdon1996geographically, fotheringham2017multiscale, comber2024multiscale]. These approaches have illuminated how environmental relationships vary across elevation gradients, climate zones, and land-use types, thereby supporting targeted interventions where specific drivers exert the greatest influence [gelfand2003spatial, finley2011comparing]. Among popular SVC methods, geographically weighted regression (GWR) [brunsdon1996geographically] computes local coefficients via distance-weighted kernels; multiscale GWR (MGWR) [fotheringham2017multiscale] extends this by allowing each predictor to operate at its own spatial scale through variable-specific bandwidths; and the geographical Gaussian process generalized additive model (GGP-GAM) [comber2024multiscale] embeds spatial smoothers within a generalized additive model framework, yielding a unified model with strong theoretical support. These methods are highly effective at capturing spatial variation when it is present. However, identifying which predictors truly have practically important spatial effects, and which have negligible effects that can be screened out, remains a distinct challenge that requires integrated spatial variable selection [wheeler2005multicollinearity, finley2011comparing]. While there is a well-developed literature on spatial variable selection in imaging [goldsmith2014smooth, smith2003assessing], in econometrics [scheel2013bayesian], in public health [choi2018bayesian] and in environmental studies [meyer2019importance], to the best of our knowledge, no prior study provides a principled spatial variable selection framework specifically aimed at identifying spatially varying environmental drivers of vegetation indices.
@@
To address this gap, we develop a Bayesian spatially varying coefficient (SVC) regression model for EVI that accommodates a broad set of environmental drivers and enables location-specific identification of important versus unimportant predictors. Each predictor-specific coefficient surface is expressed using a tensor product B-spline basis, and the associated basis coefficient vector is endowed with a Bayesian group lasso prior [yuan2006model, xu2015bayesian]. This prior induces predictor-level shrinkage: predictors with negligible influence are encouraged to have coefficient surfaces pulled toward zero, while those with practically important effects retain nontrivial spatial patterns. Posterior inference is carried out via Markov chain Monte Carlo, yielding full posterior distributions and pointwise 95% credible intervals for each coefficient surface. For each predictor, we then construct a spatial significance map by determining, at every observed location, whether the 95% credible interval includes zero. This provides a spatially explicit characterization of predictor importance for EVI over the entire domain. To summarize importance at the regional scale, we introduce a spatial coverage probability (SCP) for each predictor, defined as the proportion of locations where its coefficient is credibly nonzero. Together, the spatial significance maps and SCP summaries yield spatially-varying significance of each predictor to predict EVI in our MODIS-based vegetation analysis of the western United States.
@@
Application to MODIS vegetation data demonstrates the practical value of our framework. We find that near-infrared reflectance, red reflectance, and gross primary productivity exhibit spatially extensive, credibly nonzero effects across diverse environmental gradients, marking them as primary, large-scale drivers of vegetation condition. In contrast, mid-infrared reflectance and land cover type have coefficients that are not credibly different from zero over most of the domain, indicating that their added explanatory power for EVI is limited in this setting. These findings are consistent with established vegetation science, where NIR and red bands form the basis of most greenness indices and are tightly linked to photosynthetic activity and productivity [e.g., tucker1979red, huete2002overview, running2004continuous, pettorelli2005using]. Predictors with widespread, credibly nonzero effects warrant focused scientific interpretation and ongoing monitoring, whereas predictors with negligible influence can be omitted or modeled more parsimoniously without loss of interpretability or predictive ability.
@@ 2.1 Model Specification
Ecological relationships often exhibit spatial heterogeneity driven by environmental gradients, resource availability, and biophysical constraints [turner2005landscape, pickett1995landscape]. Spatially varying coefficient (SVC) models provide a framework for investigating such heterogeneity by allowing regression coefficients to vary over space. The model takes the form
@@
where $y({\boldsymbol{z}}_{i})$ and ${\boldsymbol{x}}({\boldsymbol{z}}_{i})=(x_{1}({\boldsymbol{z}}_{i}),...,x_{m}({\boldsymbol{z}}_{i}))^{T}$ represent the response and predictor variables, ${\boldsymbol{z}}_{i}=(u_{i},v_{i})$ represents spatial coordinates and $\beta_{j}({\boldsymbol{z}}_{i})$ are local regression coefficients. The idiosyncratic errors are taken to be independent and identically distributed (i.i.d.), $\varepsilon_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2})$. The SVC employs a separate regression for each spatial location, with these local modeling approaches capturing how the relationship between vegetation indices and environmental drivers varies across the landscape. A standard Bayesian implementation specifies Gaussian process (GP) priors for the coefficient surfaces $\beta_{j}({\boldsymbol{z}})$ [guhaniyogi2022distributed, guhaniyogi2023distributed]. While GP-based SVC models are effective for describing smoothly varying effects, they are less suitable for our MODIS vegetation application, where a central goal is to distinguish environmental drivers with substantively nonzero effects from those with negligible influence that should be screened out.
@@
where $\{\psi_{l}(\boldsymbol{z})\}_{l=1}^{L}$ is a chosen set of spatial basis functions and $\alpha_{jl}$ are the corresponding basis coefficients. A wide range of basis systems have been proposed for spatial regression, including B-splines, radial basis functions, and wavelets [e.g., eilers1996flexible, wahba1990spline, kammann2003geoadditive, ruppert2003semiparametric]. For our empirical analysis, we use tensor-product B-spline bases to flexibly approximate spatial coefficient surfaces while enabling shrinkage and selection at the predictor level, described in the next section.
@@ 2.2 Hierarchical Prior with Group lasso
This hierarchical structure induces a group lasso penalty. Integrating out the group-specific variance parameters $\tau_{j}^{2}$ produces a prior on $\boldsymbol{\alpha}_{j}$ that penalizes the $\ell_{2}$ norm $\|\boldsymbol{\alpha}_{j}\|_{2}=\sqrt{\sum_{l=1}^{L}\alpha_{jl}^{2}}$ [yuan2006model]. This group penalty shrinks entire coefficient vectors toward zero simultaneously, unlike standard lasso penalties that shrink individual coefficients independently. When applied to spatial basis coefficients, the prior encourages setting entire spatial surfaces $\beta_{j}({\boldsymbol{z}})$ to zero rather than reducing magnitudes at individual locations.
@@
This hierarchical structure differs from standard Bayesian group lasso applications [xu2015bayesian] where groups correspond to sets of related predictors or predefined variable clusters. In our spatial context, each group comprises the $L$ basis function coefficients $\boldsymbol{\alpha}_{j}$ that jointly define the $j$th predictor’s spatial surface. This spatial grouping structure enables the prior to distinguish between predictors with substantively nonzero coefficient surfaces and predictors with negligible influence on the response.
@@ 2.3 Posterior Inference: Spatial Significance Map and Spatial Coverage Probability
The resulting MCMC algorithm alternates between these conditional updates until convergence, see Algorithm 1 for details. We assess convergence using four independent chains with the Gelman-Rubin diagnostic $\hat{R}$ [gelman1992inference], where values below 1.1 indicate satisfactory convergence, along with visual inspection of trace plots for proper mixing.
@@
This quantity is conceptually different from standard posterior inclusion probabilities in Bayesian variable selection [george1993variable, barbieri2004optimal], which assign a single binary inclusion indicator to each predictor. SCP instead offers a summarization of the spatial significance map, which describes where in space a predictor matters. We deem a predictor “spatially informative” if $\text{SCP}_{j}>0.5$, meaning it has a credibly nonzero effect in more than half of the grid locations.
@@ 3.1 Simulation Design
We conduct simulation studies to evaluate the performance of the proposed Bayesian spatial group lasso (BSGL) and compare it with two established SVC methods: GGP-GAM [comber2024multiscale] and MGWR [fotheringham2017multiscale]. Simulations are conducted across various configurations with sample sizes $n\in\{1000,2000,5000,10000\}$ and number of predictors $m\in\{5,7,10\}$. For each configuration, spatial coordinates ${\boldsymbol{z}}_{i}=(u_{i},v_{i})$ are uniformly sampled from the domain $0,20]\times[0,20]$. Following the simulation design of [comber2024multiscale], predictor variables $x_{j}({\boldsymbol{z}}_{i})$ are generated from standardized normal distributions and then rescaled to the interval $[0,1]$. The idiosyncratic errors are simulated following $\varepsilon_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,\sigma^{2})$ with $\sigma^{2}=0.1$.
@@
We employ B-spline basis functions to represent the spatial coefficients. Following bai2019fast, the number of basis functions $L$ is selected from a grid consisting of $\{16,25,36,49\}$, corresponding to $4\times 4$, $5\times 5$, $6\times 6$, and $7\times 7$ tensor product B-spline bases with equal resolution in both spatial dimensions. The choice of optimal $L$ is made through 5-fold cross-validation. Hyperparameters for the group lasso prior $\lambda^{2}\sim\text{Ga}(a_{\lambda},b_{\lambda})$ are also selected via cross-validation [wood2016smoothing] from a grid where $a_{\lambda}\in\{15,30,35,40,45\}$ and $b_{\lambda}\in\{0.01,0.1,1\}$. Cross-validation evaluates all combinations using 1000 MCMC iterations with 200 warm-up samples for computational efficiency, selecting the configuration that minimizes the average mean squared prediction error $\text{MSPE}=\frac{1}{n_{I}}\sum_{i\in I}(y({\boldsymbol{z}}_{i})-\widehat{y({\boldsymbol{z}}_{i})})^{2}$, where $I$ denotes the test set indices and $\widehat{y({\boldsymbol{z}}_{i})}$ is the predicted response. Section 2 in the supplementary material offers more discussion on this.
The final model with the optimal choices of $L,a_{\lambda},b_{\lambda}$ are fitted using four independent Markov chains to ensure reliable inference. Each chain runs for 5000 iterations with the first 500 iterations discarded as warm-up, yielding 4,500 posterior samples for inference. Convergence is assessed using the Gelman-Rubin statistic $\hat{R}$ [gelman1992inference], effective sample size, and visual inspection of trace plots. In all simulation scenarios, we achieved $\hat{R}<1.01$ for all parameters, and trace plots for key hyperparameters in a representative simulation (see Section 3 of the supplementary material) demonstrate good mixing across all chains.

### v2

@@ Abstract
Keywords: Spatially varying coefficients, Bayesian variable selection, group lasso, spatial statistics, vegetation indices
@@ 1 Introduction
The western United States is undergoing rapid and unprecedented ecological change, with vegetation increasingly exposed to severe drought stress and heightened wildfire risk (Abatzoglou and Williams, 2016; Williams et al., 2019, 2020). These changes carry enormous economic costs: wildfires in the region are estimated to cost the U.S. economy \$394–\$893 billion annually, with damages exceeding \$90 billion during 2017–2021 alone. Accurately identifying which environmental factors govern vegetation health and productivity is therefore crucial for predicting ecosystem responses and designing effective adaptation strategies (Running et al., 2004; Pettorelli et al., 2005).
@@
This challenge is amplified by the region’s extreme environmental heterogeneity. The western United States spans from coastal rainforests receiving over 3000 mm of annual precipitation to interior deserts with less than 200 mm (Diffenbaugh et al., 2015), leading to fundamentally different ecological constraints. At high elevations, temperature often limits productivity, whereas in arid lowlands, moisture availability is the dominant control (Sun et al., 2024; Mukhtar et al., 2025). Effective vegetation monitoring in such a setting requires identifying which environmental variables exert substantively nonzero effects on vegetation across space, and determining where within the domain those effects are credibly nonzero.
@@
We address this problem using a predictive modeling framework for MODIS Enhanced Vegetation Index (EVI) data from the western United States, leveraging a rich suite of environmental predictors: spectral reflectance bands (red, near-infrared, blue, mid-infrared), ecosystem productivity and energy fluxes (Gross Primary Productivity, Latent Heat Flux), observation geometry (view zenith angle, sun zenith angle, relative azimuth angle), and land surface characteristics (land cover type). Standard linear regression models assume that coefficients are spatially constant, potentially obscuring important geographic variation in these relationships (Fotheringham et al., 2002). At the same time, not all candidate drivers have effects that are credibly different from zero across the region. Systematically screening predictors for non-negligible effects improves interpretability and directs scientific focus toward drivers that truly matter. Predictors with spatially extensive, credibly nonzero effects motivate spatially targeted interpretation and monitoring efforts, whereas predictors with negligible influence can often be excluded or modeled more simply, reducing monitoring and modeling complexity without sacrificing predictive accuracy.
@@
Spatially varying coefficient (SVC) models are well suited to this problem because they allow regression relationships to change smoothly across space, producing coefficient surfaces that capture geographic variation in environmental effects (Brunsdon et al., 1996; Fotheringham et al., 2017; Comber et al., 2024). These approaches have illuminated how environmental relationships vary across elevation gradients, climate zones, and land-use types, thereby supporting targeted interventions where specific drivers exert the greatest influence (Gelfand et al., 2003; Finley, 2011). Among popular SVC methods, geographically weighted regression (GWR) (Brunsdon et al., 1996) computes local coefficients via distance-weighted kernels; multiscale GWR (MGWR) (Fotheringham et al., 2017) extends this by allowing each predictor to operate at its own spatial scale through variable-specific bandwidths; and the geographical Gaussian process generalized additive model (GGP-GAM) (Comber et al., 2024) embeds spatial smoothers within a generalized additive model framework, yielding a unified model with strong theoretical support. These methods are highly effective at capturing spatial variation when it is present. However, identifying which predictors truly have practically important spatial effects, and which have negligible effects that can be screened out, remains a distinct challenge that requires integrated spatial variable selection (Wheeler and Tiefelsdorf, 2005; Finley, 2011). While there is a well-developed literature on spatial variable selection in imaging (Goldsmith et al., 2014; Smith et al., 2003), in econometrics (Scheel et al., 2013), in public health (Choi and Lawson, 2018) and in environmental studies (Meyer et al., 2019), to the best of our knowledge, no prior study provides a principled spatial variable selection framework specifically aimed at identifying spatially varying environmental drivers of vegetation indices.
@@
To address this gap, we develop a Bayesian spatially varying coefficient (SVC) regression model for EVI that accommodates a broad set of environmental drivers and enables location-specific identification of important versus unimportant predictors. Each predictor-specific coefficient surface is expressed using a tensor product B-spline basis, and the associated basis coefficient vector is endowed with a Bayesian group lasso prior (Yuan and Lin, 2006; Xu and Ghosh, 2015). This prior induces predictor-level shrinkage: predictors with negligible influence are encouraged to have coefficient surfaces pulled toward zero, while those with practically important effects retain nontrivial spatial patterns. Posterior inference is carried out via Markov chain Monte Carlo, yielding full posterior distributions and pointwise 95% credible intervals for each coefficient surface. For each predictor, we then construct a spatial significance map by determining, at every observed location, whether the 95% credible interval includes zero. This provides a spatially explicit characterization of predictor importance for EVI over the entire domain. To summarize importance at the regional scale, we introduce a spatial coverage probability (SCP) for each predictor, defined as the proportion of locations where its coefficient is credibly nonzero. Together, the spatial significance maps and SCP summaries yield spatially-varying significance of each predictor to predict EVI in our MODIS-based vegetation analysis of the western United States.
@@
Application to MODIS vegetation data demonstrates the practical value of our framework. We find that near-infrared reflectance, red reflectance, and gross primary productivity exhibit spatially extensive, credibly nonzero effects across diverse environmental gradients, marking them as primary, large-scale drivers of vegetation condition. In contrast, mid-infrared reflectance and land cover type have coefficients that are not credibly different from zero over most of the domain, indicating that their added explanatory power for EVI is limited in this setting. These findings are consistent with established vegetation science, where NIR and red bands form the basis of most greenness indices and are tightly linked to photosynthetic activity and productivity (e.g., Tucker, 1979; Huete et al., 2002; Running et al., 2004; Pettorelli et al., 2005). Predictors with widespread, credibly nonzero effects warrant focused scientific interpretation and ongoing monitoring, whereas predictors with negligible influence can be omitted or modeled more parsimoniously without loss of interpretability or predictive ability.
@@ 2.1 Model Specification
Ecological relationships often exhibit spatial heterogeneity driven by environmental gradients, resource availability, and biophysical constraints (Turner, 2005; Pickett and Cadenasso, 1995). Spatially varying coefficient (SVC) models provide a framework for investigating such heterogeneity by allowing regression coefficients to vary over space. The model takes the form
@@
where $y({\boldsymbol{z}}_{i})$ and ${\boldsymbol{x}}({\boldsymbol{z}}_{i})=(x_{1}({\boldsymbol{z}}_{i}),...,x_{m}({\boldsymbol{z}}_{i}))^{T}$ represent the response and predictor variables, ${\boldsymbol{z}}_{i}=(u_{i},v_{i})$ represents spatial coordinates and $\beta_{j}({\boldsymbol{z}}_{i})$ are local regression coefficients. The idiosyncratic errors are taken to be independent and identically distributed (i.i.d.), $\varepsilon_{i}\stackrel{{\scriptstyle\text{i.i.d.}}}{{\sim}}N(0,\sigma^{2})$. The SVC employs a separate regression for each spatial location, with these local modeling approaches capturing how the relationship between vegetation indices and environmental drivers varies across the landscape. A standard Bayesian implementation specifies Gaussian process (GP) priors for the coefficient surfaces $\beta_{j}({\boldsymbol{z}})$ (Guhaniyogi et al., 2022, 2023). While GP-based SVC models are effective for describing smoothly varying effects, they are less suitable for our MODIS vegetation application, where a central goal is to distinguish environmental drivers with substantively nonzero effects from those with negligible influence that should be screened out.
@@
where $\{\psi_{l}(\boldsymbol{z})\}_{l=1}^{L}$ is a chosen set of spatial basis functions and $\alpha_{jl}$ are the corresponding basis coefficients. A wide range of basis systems have been proposed for spatial regression, including B-splines, radial basis functions, and wavelets (e.g., Eilers and Marx, 1996; Wahba, 1990; Kammann and Wand, 2003; Ruppert et al., 2003). For our empirical analysis, we use tensor-product B-spline bases to flexibly approximate spatial coefficient surfaces while enabling shrinkage and selection at the predictor level, described in the next section.
@@ 2.2 Hierarchical Prior with Group lasso
This hierarchical structure induces a group lasso penalty. Integrating out the group-specific variance parameters $\tau_{j}^{2}$ produces a prior on $\boldsymbol{\alpha}_{j}$ that penalizes the $\ell_{2}$ norm $\|\boldsymbol{\alpha}_{j}\|_{2}=\sqrt{\sum_{l=1}^{L}\alpha_{jl}^{2}}$ (Yuan and Lin, 2006). This group penalty shrinks entire coefficient vectors toward zero simultaneously, unlike standard lasso penalties that shrink individual coefficients independently. When applied to spatial basis coefficients, the prior encourages setting entire spatial surfaces $\beta_{j}({\boldsymbol{z}})$ to zero rather than reducing magnitudes at individual locations.
@@
This hierarchical structure differs from standard Bayesian group lasso applications (Xu and Ghosh, 2015) where groups correspond to sets of related predictors or predefined variable clusters. In our spatial context, each group comprises the $L$ basis function coefficients $\boldsymbol{\alpha}_{j}$ that jointly define the $j$th predictor’s spatial surface. This spatial grouping structure enables the prior to distinguish between predictors with substantively nonzero coefficient surfaces and predictors with negligible influence on the response.
@@ 2.3 Posterior Inference: Spatial Significance Map and Spatial Coverage Probability
The resulting MCMC algorithm alternates between these conditional updates until convergence, see Algorithm 1 for details. We assess convergence using four independent chains with the Gelman-Rubin diagnostic $\hat{R}$ (Gelman and Rubin, 1992), where values below 1.1 indicate satisfactory convergence, along with visual inspection of trace plots for proper mixing.
@@
This quantity is conceptually different from standard posterior inclusion probabilities in Bayesian variable selection (George and McCulloch, 1993; Barbieri and Berger, 2004), which assign a single binary inclusion indicator to each predictor. SCP instead offers a summarization of the spatial significance map, which describes where in space a predictor matters. We deem a predictor “spatially informative” if $\text{SCP}_{j}>0.5$, meaning it has a credibly nonzero effect in more than half of the grid locations.
@@ 3.1 Simulation Design
We conduct simulation studies to evaluate the performance of the proposed Bayesian spatial group lasso (BSGL) and compare it with two established SVC methods: GGP-GAM (Comber et al., 2024) and MGWR (Fotheringham et al., 2017). Simulations are conducted across various configurations with sample sizes $n\in\{1000,2000,5000,10000\}$ and number of predictors $m\in\{5,7,10\}$. For each configuration, spatial coordinates ${\boldsymbol{z}}_{i}=(u_{i},v_{i})$ are uniformly sampled from the domain $0,20]\times[0,20]$. Following the simulation design of Comber et al. (2024), predictor variables $x_{j}({\boldsymbol{z}}_{i})$ are generated from standardized normal distributions and then rescaled to the interval $[0,1]$. The idiosyncratic errors are simulated following $\varepsilon_{i}\stackrel{{\scriptstyle\text{iid}}}{{\sim}}N(0,\sigma^{2})$ with $\sigma^{2}=0.1$.
@@
We employ B-spline basis functions to represent the spatial coefficients. Following Bai et al. (2019), the number of basis functions $L$ is selected from a grid consisting of $\{16,25,36,49\}$, corresponding to $4\times 4$, $5\times 5$, $6\times 6$, and $7\times 7$ tensor product B-spline bases with equal resolution in both spatial dimensions. The choice of optimal $L$ is made through 5-fold cross-validation. Hyperparameters for the group lasso prior $\lambda^{2}\sim\text{Ga}(a_{\lambda},b_{\lambda})$ are also selected via cross-validation (Wood et al., 2016) from a grid where $a_{\lambda}\in\{15,30,35,40,45\}$ and $b_{\lambda}\in\{0.01,0.1,1\}$. Cross-validation evaluates all combinations using 1000 MCMC iterations with 200 warm-up samples for computational efficiency, selecting the configuration that minimizes the average mean squared prediction error $\text{MSPE}=\frac{1}{n_{I}}\sum_{i\in I}(y({\boldsymbol{z}}_{i})-\widehat{y({\boldsymbol{z}}_{i})})^{2}$, where $I$ denotes the test set indices and $\widehat{y({\boldsymbol{z}}_{i})}$ is the predicted response. Section 2 in the supplementary material offers more discussion on this.
The final model with the optimal choices of $L,a_{\lambda},b_{\lambda}$ are fitted using four independent Markov chains to ensure reliable inference. Each chain runs for 5000 iterations with the first 500 iterations discarded as warm-up, yielding 4,500 posterior samples for inference. Convergence is assessed using the Gelman-Rubin statistic $\hat{R}$ (Gelman and Rubin, 1992), effective sample size, and visual inspection of trace plots. In all simulation scenarios, we achieved $\hat{R}<1.01$ for all parameters, and trace plots for key hyperparameters in a representative simulation (see Section 3 of the supplementary material) demonstrate good mixing across all chains.